# for Unseen Language Generalization

### v2

@@ Abstract
Pretrained language models (PLMs) have become remarkably adept at task and language generalization. Nonetheless, they often fail dramatically when faced with unseen languages, posing a significant problem for diversity and equal access to PLM technology. In this work, we present LinguAlchemy, a regularization technique that incorporates various aspects of languages covering typological, geographical, and phylogenetic constraining the resulting representation of PLMs to better characterize the corresponding linguistics constraints. LinguAlchemy significantly improves the accuracy performance of mBERT and XLM-R on unseen languages by $\sim$18and $\sim$2%, respectively compared to fully fine-tuned models and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search. LinguAlchemy enables better cross-lingual generalization to unseen languages which is vital for better inclusivity and accessibility of PLMs.
@@ 1 Introduction
Significant advancements in language processing technology have been achieved through the development of PLMs, leading to a commendable proficiency in language comprehension and generation Devlin et al. (2019); Liu et al. (2019); Conneau et al. (2020); Sanh et al. (2022); Lewis et al. (2019); Raffel et al. (2023); Li et al. (2021); Cahyawijaya et al. (2021); Wilie et al. (2020). However, there remains a notable deficiency in the ability of these models to generalize effectively to unseen languages, resulting in a considerable performance reduction of PLMs across thousands of unseen languages. To mitigate this problem, efforts to develop efficient language adaptation approaches are underway, focusing on the incorporation of these unseen languages to PLMs Pfeiffer et al. (2021b); Alabi et al. (2022); Ebrahimi et al. (2022); Goyal et al. (2021).
@@
Incorporating new unseen languages has been a longstanding problem in natural language processing (NLP) research, especially given that most of these unseen languages are low-resource and underrepresented making PLMs difficult to adapt to these languages. MAD-X Pfeiffer et al. (2020b) employs a language adapter to learn new unseen languages by incorporating language adapters that mitigate the risk of forgetting pre-trained knowledge which is known as the curse-of-mulitlinguality. Nonetheless, this approach requires training for generalizing to new unseen languages which makes it costly and difficult to scale to thousands of languages. MAD-G Ansell et al. (2021) and Udapter Üstün et al. (2020) further generalize this approach by utilizing a linguistic-driven contextual parameter generator (CPG) module to generate language-specific parameters allowing the models to generalize to other languages with similar linguistic characteristics. Recently, Rathore et al. (2023) introduced ZGUL, which combines representations over multiple language adapters to generate the unseen language representation. Despite the effectiveness, all these approaches largely rely on two assumptions, i.e., 1) strict categorization of languages and 2) knowing the language category of the query apriori. The first assumption disregards the fact that linguistic phenomena such as code-mixing may occur in the query. While the second assumption might cause performance degradation due to the error propagation from the language identification module Adilazuarda et al. (2023).
To overcome those limitations, in this work, we introduce LinguAlchemy. Unlike adapter-based approaches which isolate the capability to understand different languages over multiple language-specific adapters, LinguAlchemy learns a shared representation across multiple languages, allowing the model to leverage shared knowledge between languages. Such behavior is attainable by eliminating the language-specific module from the model and instead utilizing a regularization to learn language-specific knowledge. LinguAlchemy allows the model to perform inference without knowing the language of the query prior. Our evaluation suggests that LinguAlchemy can improve mBERT and XLM-R generalization in unseen languages while maintaining the performance on the high-resource languages without requiring information on the language of the query apriori.
@@
    We propose LinguAlchemy, a regularization method that improves unseen language performance on language models and aligns them to arbitrary languages.
@@
    We demonstrate strong performance on unseen languages for models trained with LinguAlchemy.
@@
    We introduced a dynamic scaling method to scale the classification and auxiliary loss factors used in the fine-tuning stage.
2 Related Works
@@
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation. These models excel in abstract linguistic generalization by capturing complex linguistic patterns and understanding structural positions and thematic roles, which are crucial for interpreting language semantics. Research in this area Ganesh et al. (2021) has provided critical insights that enable these models to process and generate human language effectively. The studies have explored how these models grasp intricate linguistic features, including syntax and semantics, thereby enhancing their performance across a wide range of language tasks Rathore et al. (2023).
In parallel, the development of resources like URIEL vector and lang2vec utility Littell et al. (2017) has been instrumental in extending the reach of multilingual NLP, particularly for less-resourced languages. These tools provide vector representations of languages, leveraging typological, geographical, and phylogenetic data, thus offering a structured approach to understanding linguistic diversity. Complementing this, recent research has conducted a comprehensive survey on the utilization of typological information in NLP, highlighting its potential in guiding the development of multilingual NLP technologies Ponti et al. (2019). This survey emphasized the underutilization of typological features in existing databases and the need for integrating data-driven induction of typological knowledge into machine learning algorithms.
Recent advancements in prefix tuning Li and Liang (2021) and subspace learning Zhang et al. (2020) have contributed significantly to improving generalization in PLMs. These methods focus on learning prefix subspaces to stabilize the direct learning of embeddings, addressing the instability issues present in earlier approaches. Two notable methods are the MAD-X Pfeiffer et al. (2020b) and MAD-G Ansell et al. (2021) frameworks, which employ adapter-based techniques for multi-task cross-lingual transfer, highlighting modularity and parameter efficiency. However, it inherits the limitations of the pretrained multilingual models, such as the limited capacity to adapt effectively to low-resource and unseen languages. Furthermore, while the framework facilitates adaptation to specific target languages, it may bias the model towards these languages, potentially impacting its performance on other languages.
@@
However, despite these advancements, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender (2011); Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017). This situation highlights the complexity of scaling and generalizing these models effectively and underscores the need for more sophisticated approaches in model training and adaptation to ensure broader and more equitable language coverage.
@@ 3 Unseen Languages Adaptation with LinguAlchemy
3.1 Does Multilingual LMs capture Linguistic Constraint?
In this work, we define the linguistic knowledge as a vector gathered from URIEL vector (Littell et al., 2017). We chose three distinct linguistic knowledge from the database, namely ’syntax_knn’, ’syntax_average’ [^1] In this work, we chose the ’knn’ and ’average syntax features. These include consensus values (like averages) and predicted values (such as kNN regressions based on phylogenetic or geographical neighbors), and ’geo’ features. The choice of ’syntax_knn’ and ’syntax_average’ is motivated by the typological nature of syntax. Syntax in languages varies widely; hence, by using aggregate measures like averages and k-nearest neighbors (kNN), we can capture a more general representation of syntactic features across languages. These features include consensus values, like averages, and predicted values, such as kNN regressions based on phylogenetic or geographical neighbors.
We excluded phonological features, as our focus is not on speech but on textual data. Phonology primarily pertains to spoken languages and would not add significant value to our analysis of written language structures. Additionally, we excluded language family features (’family’). These features are typically binary, indicating whether a language belongs to a particular subfamily. While they can be useful for language identification tasks, they tend to be sparse and may not provide the granularity needed for our study. By focusing on syntax and geographical features, our approach aims to encompass a broader and more nuanced understanding of linguistic variations in a multilingual context.
Syntax Feature
These feature vectors denote a typological feature that is adapted from several sources including World Atlas of Language Structures (WALS), Syntactic Structures of World Languages Collins (2010), and short prose descriptions on typological features in Ethnologue Lewis (2009). Syntax vectors captures information about the syntactic properties of languages which derived from large-scale typological databases, which document the structural and semantic variation across different languages. These syntax features in the URIEL vector are utilized to represent languages in vector form, allowing for the analysis and comparison of languages based on their syntactic properties.
Geographical Feature
On the other hand, geographical features represent languages in terms of their geographical properties. The inclusion of ’geo’ features aims to capture geographical attributes of languages. Geographic factors can significantly influence language evolution and structure, making them crucial for understanding linguistic variations. This feature expresses geographical location with a fixed number of dimensions that each represents the “great circle” distance—from the language in question to a fixed point on the Earth’s surface. By incorporating geographical information into language vectors, URIEL and lang2vec provide a more comprehensive view of languages, considering not only their structural and semantic properties but also their geographical context.
@@ 3.2 Proof of Concept
To explore whether multilingual language models (MLMs), such as Multilingual BERT (mBERT), capture the linguistic constraints as defined by the URIEL vectors, we align the mBERT language embeddings to the linguistic knowledge vectors. This projection is quantitatively measured to assess the representation of both seen and unseen languages within the model. The fundamental question we address is the extent to which mBERT’s embeddings correspond to the typological and geographical features encapsulated in the URIEL vectors.
@@
Figure 2 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot accentuates the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization underscores the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers prompts a closer examination of mBERT’s coverage and consistency in representing diverse linguistic features. We also formally defined the language representation alignment in the Appendix B.
@@
We introduce LinguAlchemy as an approach that intuitively aligns model representations with linguistic knowledge, leveraging URIEL vectors. This is operationalized through an auxiliary loss function, involving the training process with a nuanced understanding of linguistic characteristics.
In LinguAlchemy, we enhance the fine-tuning of encoder models such as mBERT for downstream tasks by not only using the regular classification loss but also introducing a novel linguistic regularization term. This is achieved through the implementation of a URIEL loss, designed to align the model’s representations with linguistic knowledge derived from URIEL vectors. Specifically, this process involves applying a linear projection to the model’s pooled output, which aligns it with the URIEL vector space. The URIEL loss is quantified as the mean squared error (MSE) between the projected model outputs and the corresponding URIEL vectors. This dual approach, combining classification loss and URIEL loss, allows for a more linguistically informed model training, enhancing the model’s ability to capture and reflect complex linguistic patterns and relationships.
$L_{uriel}(R,U)=\frac{1}{N}\sum_{i=1}^{N}\|R_{i}-U_{i}\|^{2}$
where $R$ represents the model-generated representations, $U$ denotes the URIEL vectors, and $N$ is the number of data points. To generate the model representation, we take the output representation from the $CLS$ token and multiply it with a new, trainable projection layer to transform the vector size so that they are compatible.
Note that there may be discrepancies between the scales of the standard classification loss and the URIEL loss. To address this, we introduce an optional hyperparameter, denoted as $\lambda$, to scale the URIEL loss appropriately.
@@
$L=\lambda_{cls}*L_{cls}+\lambda_{uriel}*L_{uriel}(R,U)$
@@
    AlchemyScale: This method dynamically adjusts the scaling factor $\lambda$ during training. It initiates with scaling factors set relative to the mean of initial losses, ensuring proportional importance to each loss component. Subsequently, these factors are updated periodically using an Exponential Moving Average (EMA) method, with a conservative adjustment for specific loss components to maintain stability. This approach ensures an optimal balance between different loss components, adjusting the contribution of each to the total loss based on predefined criteria.
@@
    AlchemyTune: Here, $\lambda$ is conceptualized as a trainable parameter within the model’s architecture. Initialized as part of the model’s parameters, $\lambda$ undergoes optimization during the training process. This method applies the scaling factors to loss components, and an additional mini_loss—representing the deviation of the sum of scaling factors from unity—is computed. This mini_loss is used for backpropagation, enabling the scaling factors to adapt based on the training dynamics and dataset specifics.
@@ 4 Experiment Setting
In our experiments, we utilize the MASSIVE Xu et al. (2022), which is a comprehensive collection of multilingual data incorporating intent classification tasks. This dataset is particularly notable for its inclusion of a diverse range of languages, with various language families, genera, and scripts. Specifically, we split MASSIVE into 25 languages that are ’seen’ during finetuning and the rest 27 languages that are ’unseen’, which we exclusively used for evaluation. This splitting is based on the language adapters availability as outlined in the prior research of (Pfeiffer et al., 2020a), which we utilized in the AdapterFusion experiment for our baseline model. For a detailed breakdown of the languages used, including their respective families, genera, and script can be found in Appendix A.
@@
Our research employs two state-of-the-art language models: Multilingual BERT Base (mBERTBase) and XLM-RoBERTa Base (XLM-RBase). These models are chosen for their robust performance across a wide range of multilingual NLP tasks, making them ideal for our intent classification and language identification objectives.
Hyperparameters
The following hyperparameters are used in our training process:
  • Learning rate: $5\times 10^{-5}$
  • Training epochs: 30
  • Performance metric: Accuracy
  • Early stopping criterion: 3 epochs without improvement

### v3

@@ Abstract
Pretrained language models (PLMs) have become remarkably adept at task and language generalization. Nonetheless, they often fail when faced with unseen languages. In this work, we present LinguAlchemy, a regularization method that incorporates various linguistic information covering typological, geographical, and phylogenetic features to align PLMs representation to the corresponding linguistic information on each language. Our LinguAlchemy significantly improves the performance of mBERT and XLM-R on low-resource languages in multiple downstream tasks such as intent classification, news classification, and semantic relatedness compared to fully finetuned models and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search.
@@ 1 Introduction
Significant advancements in language processing technology have been achieved through the development of PLMs, leading to a commendable proficiency in language comprehension and generation Devlin et al. (2019); Liu et al. (2019); Lewis et al. (2019); Li et al. (2021); Sanh et al. (2022); Raffel et al. (2023). The advancement has been further expanded to non-English languages Conneau et al. (2020); Martin et al. (2020); Wilie et al. (2020); Kakwani et al. (2020); Cahyawijaya et al. (2021, 2024). However, there is still a gap in these models’ ability to generalize effectively to unseen languages. To mitigate this problem, there are some efforts to develop efficient language adaptation methods to unseen languages Pfeiffer et al. (2021b); Goyal et al. (2021); Alabi et al. (2022); Ebrahimi et al. (2022); Yong et al. (2023).
@@
Previous approaches Rathore et al. (2023); Üstün et al. (2022); Pfeiffer et al. (2020b); Ansell et al. (2021) are based on the assumption that disregards the fact that in a real-world application, there is usually no language information from the user, highlighting the importance of the multilingual robustness of a language model. The second assumption might cause performance degradation due to the error propagation from the language identification module Adilazuarda et al. (2023). However, these methods inherit the limitations of the pretrained multilingual models, such as the limited capacity to adapt effectively to low-resource and unseen languages. Furthermore, while the framework facilitates adaptation to specific target languages, it may bias the model towards them, potentially impacting its performance on other languages.
In this work, we introduce LinguAlchemy, a novel method that incorporates a unified representation across multiple languages to enable the model for utilizing the shared linguistic knowledge. Our approach differs from adapter-based approaches which often segment language understanding into multiple, isolated language-specific modules. Instead, we developed a regularization technique that utilizes linguistic information directly into the model’s architecture, allowing for language-agnostic inference. Our evaluations demonstrate that LinguAlchemy not only enhances generalization capabilities of mBERT Devlin et al. (2018) and XLM-R Conneau et al. (2020) on unseen languages but also upholds robust performance across high-resource languages, all without prior knowledge of the query’s language.
Our strategy aims to refine cross-lingual generalization by leveraging linguistic features encapsulated in URIEL vectors. We hypothesize that languages with similar syntactic and typological characteristics can benefit from shared representational frameworks, significantly boosting performance in multilingual settings. This approach is particularly beneficial in contexts where language resources are limited.
@@
    We propose LinguAlchemy, a regularization method that utilizes geographical and syntactic information to foster the models’ unified representation.
@@
    LinguAlchemy is a new regularization approach that does not require any architectural change and can be adapted to different tasks and models.
@@
    We demonstrate strong performance on 50+ languages across three diverse datasets and tasks (intent classification, news classification, and semantic relatedness) for models trained with LinguAlchemy, including the languages that are not seen during pretraining.
  4. 4.
@@
    We introduce two automatic hyperparameter search methods to scale the classification and auxiliary loss factors used in the fine-tuning stage, namely dynamiclearn and dynamicscale.
2 Related Work
@@
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation. Ganesh et al. (2021) have explored how these models learn intricate linguistic features, including syntax and semantics to enhance their performance across a wide range of language tasks Rathore et al. (2023).
Incorporating new unseen languages has been a longstanding problem in the multilingual research, MAD-X Pfeiffer et al. (2020b) employ a language adapter to learn new unseen languages using language adapters that mitigate the risk of forgetting pre-trained knowledge, which is known as the curse-of-multilinguality Conneau et al. (2020). Nonetheless, this approach requires training for generalizing to new unseen languages, which makes it costly and difficult to scale to thousands of languages. MAD-G Ansell et al. (2021) and Udapter Üstün et al. (2020) further generalize this approach by utilizing a linguistic-driven contextual parameter generator (CPG) module to generate language-specific parameters, allowing the models to generalize to other languages with similar linguistic characteristics. Recently, Rathore et al. (2023) introduce ZGUL, which combines representations over multiple language adapters to generate the unseen language representation. Despite the effectiveness, all these approaches rely on two assumptions: (1) strict categorization of languages and (2) knowing the language category of the query apriori—our definition of “a priori categorization" as incorporating language-specific information into the model.
In parallel, the development of linguistically-driven resources such as the URIEL vector and the lang2vec utility Littell et al. (2017) has been notable in extending multilingual NLP research, particularly for less-resourced languages. Complementing this, Ponti et al. (2019) emphasized the underutilization of typological features in existing databases and the need for integrating data-driven approaches of typological knowledge into language models. Previous studies have focused on extending multilingual NLP using URIEL vectors Lauscher et al. (2020); Lin et al. (2019); Tan et al. (2019); Oncevay et al. (2020), but none of their approaches use URIEL vectors for alignment during the finetuning process of language models.
@@
However, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender (2011); Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017). This situation highlights the complexity of scaling and generalizing these models effectively and underscores the need for more sophisticated approaches in model training and adaptation to ensure broader and more equitable language coverage.
@@ 3 Unseen Languages Adaptation with LinguAlchemy
3.1 Does Multilingual LMs capture Linguistic Constraints?
In this work, we define the linguistic knowledge as a vector gathered from URIEL vector (Littell et al., 2017). We chose three distinct linguistic knowledge from the database, namely ’syntax_knn’, ’syntax_average’ [^1] In this work, we chose the ’knn’ and ’average syntax features. These include consensus values (like averages) and predicted values (such as kNN regressions based on phylogenetic or geographical neighbors), and ’geo’ features. The choice of ’syntax_knn’ and ’syntax_average’ is motivated by the typological nature of syntax. Syntax in languages varies widely; hence, by using aggregate measures like averages and k-nearest neighbors (kNN), we can capture a more general representation of syntactic features across languages. Note that in our experiments, we excluded phonological features and language family attributes from our analysis as they are less relevant to textual data and have limited granularity for understanding linguistic variations.
Syntax Features
These feature vectors denote a typological feature that is adapted from several sources including World Atlas of Language Structures (WALS) Dryer and Haspelmath (2013), Syntactic Structures of World Languages Collins (2010), and short prose descriptions on typological features in Ethnologue Lewis (2009).Syntax vectors capture information about the syntactic properties of languages, derived from large-scale typological databases that document the structural and semantic variation across different languages. These syntax features in URIEL are utilized to represent languages in vector form that allows the analysis and comparison of languages based on their syntactic properties.
Geographical Features
On the other hand, geographical features represent languages in terms of their geographical properties. The inclusion of ‘‘geo" features aims to capture geographical attributes of languages. This feature expresses geographical location with a fixed number of dimensions that each represents the “great circle" distance—from the language in question to a fixed point on the Earth’s surface. By incorporating geographical information into language vectors, URIEL and lang2vec provide a more comprehensive view of languages, considering not only their structural and semantic properties but also their geographical context.
@@ 3.2 Proof of Concept
We investigate whether multilingual language models (MLMs) like mBERT Devlin et al. (2018) can capture linguistic constraints by aligning mBERT language embeddings with URIEL vectors to assess how they represent seen and unseen languages. This includes examining how well mBERT’s embeddings correspond to the typological and geographical features detailed in URIEL. In Figure 2, sentence embeddings (green dots) from mBERT, derived from the last hidden state of multilingual training data, and URIEL vectors (brown dots)—structured representations from the URIEL database—are projected into the same space. A matrix $W$ is used to linearly project sentence embeddings, minimizing the mean squared error with URIEL vectors. This alignment is showcased in Figure 2 using UMAP for dimensionality reduction for visualization purpose.
@@
Figure 2 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot highlight the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization shows the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers requires us to carefully check mBERT’s coverage and consistency in representing different linguistic features. We also formally defined the language representation alignment in the Appendix B.
2 Model Feature Type Acc.(%) mBERT geo 66.14 syntax_avg+geo 65.62 syntax_avg 66.41 syntax_knn+geo 66.08 syntax_knn 66.41 syntax_knn+syntax_avg+geo 66.47 syntax_knn+syntax_avg 66.41 XLM-R geo 80.16 syntax_avg+geo 80.54 syntax_avg 80.76 syntax_knn+geo 80.48 syntax_knn 80.32 syntax_knn+syntax_avg+geo 80.80 syntax_knn+syntax_avg 80.78
Table 1: Linguistic vector ablation experiment (the highest accuracy for each model is highlighted in bold).
Table 1 shows the linguistic vector ablation experiment conducted on two models, mBERT and XLM-R. The vector ablation experiment demonstrates that the combination of syntax_knn, syntax_avg, and geographical features yields the best results. Therefore, we will use this combination of features in subsequent experiments.
@@
We introduce LinguAlchemy as an approach that intuitively aligns model representations with linguistic knowledge, leveraging URIEL vectors. This approach is applied through an auxiliary loss function that is involved in the training process with an added information of linguistic characteristics in the form of URIEL vector.
In LinguAlchemy, we enhance the fine-tuning of encoder models such as mBERT for downstream tasks by not only using the regular classification loss but also introducing a novel linguistic regularization term. This is achieved through the implementation of a URIEL loss, designed to align the model’s representations with linguistic knowledge derived from URIEL vectors. Specifically, this process involves applying a linear projection to the model’s pooled output, which aligns it with the URIEL vector space. The URIEL loss is quantified as the mean squared error (MSE) between the projected model outputs and the corresponding URIEL vectors. This dual approach allows for a more linguistically informed model training and increase the model’s ability to capture and reflect complex linguistic patterns.
$\mathcal{L}_{uriel}(Z,U)=\frac{1}{N}\sum_{i=1}^{N}\|Z_{i}-U_{i}\|^{2},$ (1)
LinguAlchemy is represented by equation 1 where $Z$ represents the model-generated representations, $U$ denotes the URIEL vectors, and $N$ is the number of data points. To generate the model representation, we take the output representation from the $CLS$ token and multiply it with a new, trainable projection layer to transform the vector size so that they are compatible.
Note that there may be discrepancies between the scales of the standard classification loss and the URIEL loss. To address this, we introduce an optional hyperparameter, denoted as $\lambda$, to scale the URIEL loss appropriately. However, finding this scaling factor requires another hyperparameter search. Therefore, we propose a new method for dynamically balancing the classification and URIEL losses using two dynamic scaling approaches.
@@
$\mathcal{L}=\lambda_{cls}*\mathcal{L}_{cls}+\lambda_{uriel}*\mathcal{L}_{uriel}(Z,U).$ (2)
@@
    AlchemyScale: This method dynamically adjusts the scaling factor $\lambda$ during training. It is initiated with scaling factors set relative to the mean of initial losses. Furthermore, these factors are updated periodically using an Exponential Moving Average (EMA) method to balance between different loss components.
@@
    AlchemyTune: Here, $\lambda$ is conceptualized as a trainable parameter within the model’s architecture. Initialized as part of the model’s parameters and optimized during the training process. This method applies the scaling factors to loss components, then an additional mini_loss is computed to represent the deviation of the sum of scaling factors.
@@ 4 Experiment Setting
In our experiments, we use MASSIVE Dataset FitzGerald et al. (2023), which is a comprehensive collection of multilingual data incorporating intent classification tasks. We split MASSIVE into 25 languages that are “seen" during finetuning and the rest 27 languages that are “unseen", which we exclusively used for evaluation. This splitting is based on the language adapters availability as outlined in the prior research of Pfeiffer et al. (2020a), which we utilized in the AdapterFusion experiment for our baseline model. For a detailed breakdown of the languages used, including their respective families, genera, and script can be found in Appendix A.
Additionally, we incorporate the MasakhaNews Dataset Adelani et al. (2023), consisting of news article classification across several African languages. This dataset tests our models against diverse journalistic styles and complex syntactic structures. For our experiments, the training languages are amh, eng, fra, hau, swa, orm, and som, while the testing languages include ibo, lin, lug, pcm, run, sna, tir, xho, and yor. Lastly, we also utilize the SemRel2024 Dataset Ousidhoum et al. (2024) for semantic relatedness task in low-resource languages. We use this dataset to evaluate LinguAlchemy’s semantic understanding and relationship extraction capabilities. We train using the languages amh, arq, ary, eng, esp, hau, kin, mar, and tel. The test set includes afr, amh, arb, arq, ary, eng, esp, hau, hin, ind, kin, and pan.
@@
Our research employs two state-of-the-art language models: Multilingual BERT Base (mBERTBase) and XLM-RoBERTa Base (XLM-RBase). In our training process, we use a learning rate of $5\times 10^{-5}$, train for 30 epochs, and measure performance based on accuracy for MASSIVE and MasakhaNews, and Pearson correlation for SemRel. Each training takes at most 5 hours using a single A100 GPU.